Karl Schmidt just scored a 96 on his last biology unit test. The mean score of the 36 students in his class was 78, and the standard deviation was 12. What is Karl’s z-score?: Karl Schmidt just scored a 96 on his last biology unit test. The mean score of the 36 students in his class was 78, and the standard deviation was 12. What is Karl’s z-score?
Answer by stanbon(18984)

Karl Schmidt just scored a 96 on his last biology unit test. The mean score of the 36 students in his class was 78, and the standard deviation was 12. What is Karl’s z-score?
---------------------
Formula:
z(x) = (x-u)/sigma
z(96) = (96-78)/12 = 18/12 = 1.5
===================================
Cheers,
Stan H.